OpenStudy (anonymous):

How do I solve: (sqrt(2x^2-11x+14))^(1/3)=2-x

OpenStudy (mertsj):

\[\sqrt{2x^2-11x+14}^\frac{1}{3}=2-x\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yup!

OpenStudy (mertsj):

So it is really: \[(2x^2-11x+14)^{\frac{1}{6}}=2-x\]

OpenStudy (anonymous):

is that 1/6?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

so would we multiple the 2-x by that too?

OpenStudy (mertsj):

The only thing I can think of to do is to raise both sides to the 6th power.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Proably use Pascal's triangle or the binomial theorem for the right side.

OpenStudy (anonymous):

wait... if we multiplied the sq.root by 2 (so as to get 1/6), wouldn't we have to multiply the right by (1/6) as well?

OpenStudy (anonymous):

wait that doesn't work... where did the 1/6 come from?

OpenStudy (mertsj):

Would you agree that: \[\sqrt{x}=x ^{\frac{1}{2}}\]

OpenStudy (mertsj):

And would you further agree that you multiply exponents when raising to a power?

OpenStudy (anonymous):

No, wouldn't it be \[\sqrt{x}=(x)^2\]

OpenStudy (mertsj):

And so if you have: \[(\sqrt{x})^{\frac{1}{3}}= [(x)^{\frac{1}{2}}]^{\frac{1}{3}}=x ^{\frac{1}{6}}\]

OpenStudy (mertsj):

So you are saying that \[\sqrt{x}=x^2\]

OpenStudy (mertsj):

??????

OpenStudy (mertsj):

So: \[\sqrt{4}=4^2=16\] ???????

OpenStudy (mertsj):

I always thought that the square root of 4 is 2 not 16

OpenStudy (anonymous):

yes, if you want to get rid of a square root, you simply square it.

OpenStudy (anonymous):

oh, you weren't trying to get RID of the square root.

OpenStudy (mertsj):

So what you are really saying is that \[(\sqrt{x})^2=x\]

OpenStudy (anonymous):

So then yes, you're right. yes, sorry.

OpenStudy (mertsj):

That's a whole different thing than what you wrote before.

OpenStudy (anonymous):

I guess it is! So could we not solve the question be first eliminating the square root?

OpenStudy (mertsj):

Actually you have two roots: a square root and a third root

OpenStudy (anonymous):

could you multiply the exponent of 1/3 by 3 (so as to cancel it) then just square the square root?

OpenStudy (mertsj):